OpenStudy (noah690):

Solve the equation: 1/3 + 2/3x = 5/6. Is the answer: x = 3/4. Thanks.

OpenStudy (anonymous):

Tell me how you got 3/4 because it is not 3/4

OpenStudy (igreen):

Yep, 3/4 is correct.

OpenStudy (igreen):

2/3 * 3/4 = 6/12 = 1/2 1/2 + 1/3 = 3/6 + 2/6 = 5/6

OpenStudy (anonymous):

Oh wait I read the question wrong sorry, Your answer actually is correct

OpenStudy (igreen):

:P

OpenStudy (noah690):

Thanks for the help.

OpenStudy (igreen):

No problem, \(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\)

OpenStudy (anonymous):

noah690 do you go to Georgia connecitons 8 grade